OpenStudy (anonymous):

Find the derivative of the function integral from 4x to 7x of (u^2-1)/(u^2+1)

OpenStudy (anonymous):

\[\int\limits_{4x}^{7x} \frac{ u^2-1 }{ u^2 + 1 }\]

OpenStudy (anonymous):

are u sure the problem like that ??

OpenStudy (anonymous):

I'm suppose to find the derivative of that integral

OpenStudy (anonymous):

yea i see.. but actually if using u's variables, so we use sign du.. but there i cant see du or dx ?

OpenStudy (anonymous):

Yeah I have no idea how to approach this problem because u-substitution wouldn't apply

OpenStudy (anonymous):

summon or tag @Hero , @hartnn ....

hero (hero):

Is it du or dx?

OpenStudy (anonymous):

du

hero (hero):

Okay

hero (hero):

Integral and Derivative are inverses of each other, so you end up with the the original function you came up with including the limits of integration: \[\left[ \frac{u^2-1}{u^2+1} \right]^{7x}_{4x}\]

hero (hero):

Then continue evaluating: \[\left[ \frac{u^2-1}{u^2+1} \right]^{7x} - \left[ \frac{u^2-1}{u^2+1} \right]_{4x}\]

hero (hero):

\[\left[ \frac{(7x)^2+1}{(7x)^2-1} \right] - \left[ \frac{(4x)^2+1}{(4x)^2-1} \right]\]

hero (hero):

I'm pretty sure you can finish that on your own.

OpenStudy (anonymous):

thank you

hero (hero):

yw